# Applications of Generative AI (GAI) for Mobile and Wireless Networking: A Survey

### v1

@@ III-D Channel Estimation
Numerous benefits can be achieved by incorporating GAI approaches into mobile networking channel estimation. First, GAI enhances the ability to detect faulty signals and provide appropriate fixes to improve channel estimation procedures. Second, GAI can streamline the resources used for channel estimation, which lowers the computational complexity and energy usage of conventional estimation techniques. Third, by learning the fundamental patterns and properties of wireless channels, GAI makes channel estimation more reliable and precise, particularly in difficult situations with little training data or time-varying channel circumstances.
@@
To verify these promising traits, the work in [97] designs a high dimensional channel estimation approach based on a deep generative network combined with a compressive sensing strategy. The experiments from [97] indicate that GAI not only uses fewer channel pilots than conventional techniques, but also improves the execution time of channel estimation procedures for one-bit quantized pilot measurements. Similarly, another study for wideband channel estimation [98] has demonstrated the efficacy of GANs in dealing with channel estimation in the conditions of limited channel pilots and low transmit SNR. Specifically, it significantly reduces the number of pilots by about 70%, even at mmWave and THz frequencies, without affecting the estimation errors. On another front, the work in [99] displays the potential of deep score-based generative models to multi-input multi-output systems, with high fidelity channel estimation for channel communication of sizes 64×256 with pilot density up to 25%.
III-E Lessons Learned and Keys Takeaway
Modern networks have undergone a revolution because of the use of GAI in network management, which has made it possible to provide accurate and high-quality data. Among GAI techniques, GANs have become popular in many applications, such as picture synthesis, text generation, and data augmentation, providing innovative and varied training data for AI models. In wireless networks, GAI is crucial for allocating resources effectively, optimizing input data to cut transmission costs and increase accuracy. Additionally, generative models have solved practical issues in channel estimation by identifying network risks with greater precision and eliminating data imbalance problems.
The landscape of network routing presents challenges, encompassing aspects like security, user verification, and data integrity. Recent research indicates that the GAI is instrumental in addressing network routing challenges. It excels in enhancing security, conducting user verification, ensuring secure communication, and fortifying against advanced attacks. Future investigations should not be confined to refining GAI algorithms for optimized mobile network operations. Instead, they should focus on fortifying data security and continually enhancing user experiences. Moreover, the dynamic potential of GAI in mobile networking requires ongoing exploration for improved creativity and security in communication networks.
In summary, the key takeaways from this section are as follows.
@@ III-E Lessons Learned and Keys Takeaway
  • GAI enhances network management by employing predictive models to optimize traffic flow and minimize downtime.
@@
  • By dynamically adjusting computing resources based on real-time demand and ensuring optimal efficiency, the resource allocation capabilities are improved.
  • Channel estimation and security benefit are achieved from generative models with improved estimation and detection accuracy.
  • GAI can strengthen network security utilizing advanced anomaly detection algorithms to identify and mitigate potential threats.
  • GAI can contribute toward data confidentiality and integrity by implementing encryption protocols and monitoring data for unauthorized access or alterations.
@@ IV Application of GAI in Wireless Security
Wireless security is a critical concern in mobile networking, with the ever-growing need to protect sensitive data and ensure secure communication. In recent years, GAI has emerged as a powerful tool with diverse applications, including anomaly detection, intrusion detection, authentication, and encryption. By harnessing the capabilities of generative models, novel approaches are being developed to enhance the resilience and effectiveness of security mechanisms in mobile networks. This section delves into the advancements, challenges, and potential directions in leveraging GAI for wireless security, contributing to the evolving landscape of mobile networking.
@@ IV-A Sensitive Data Obscurity
Data sharing has become a popular trend in intelligent systems for training machine learning models. However, this activity poses numerous risks related to user privacy, in that sensitive information may be disclosed by third parties without unauthorized access or interception [100]. In this context, GAI has recently demonstrated its promising capabilities in obfuscating sensitive data by creating synthetic data that closely mimics real datasets [101].
There are multiple methods for generating synthetic data using GAI models, such as using GAN, VAE, or GPT models. In particular, data perturbation is one of GAI’s main techniques in safely hiding sensitive data [102]. This technique uses synthetic data to complicate extraction of useful information, by applying controlled perturbations to sensitive data [103]. In addition, distributing and transforming only the sensitive elements in datasets can also increase the security of data sharing while preserving the features and statistical utility of the original data for model training activities [104]. Moreover, GAI models can apply other techniques like differential privacy and data anonymization.
GAI techniques also aid in encrypted communication. Indistinguishable noise can be created from encrypted data using generative models, complicating decryption of the original data without the proper decryption keys [105]. With this strategy, sensitive data are sent more securely over wireless networks while maintaining their confidentiality and integrity. However, despite the apparent advantages, there are issues when using GAI to obfuscate sensitive data, such as the trade-off between data value and obscurity, the scalability of generative models of huge datasets, as well as the criteria and requirements for the reliability and efficacy evaluation of data obscurity strategies.
@@ IV-B Intrusion Detection
Intrusion detection in wireless security is essential for identifying and combating threats or intrusions on mobile networks. Applying GAI techniques opens new avenues to strengthen IDSs and enhance wireless network security as a whole [106]. A valuable tool for discovering anomalies and distinguishing malicious activity in wireless network traffic is provided by GAI models like GANs. These models are capable of learning the typical patterns and behaviors of network traffic through training on large-scale datasets, allowing them to detect unforeseen attack patterns from expected norms with high accuracy [107].
Generating synthetic data samples that closely resemble actual network traffic is one of the main benefits of utilizing GAI in intrusion detection. IDS is capable of accomplishing this by enhancing its training data and producing a variety of testing and evaluation situations. IDS can increase its robustness and adaptability to various sorts of threats by creating synthetic attack patterns to improve its detection capabilities. The early discovery of zero-day attacks, vulnerabilities or exploits of which security experts are unaware, can also be helped by GAI [108]. IDS can detect suspicious actions that depart from usual behavior by continually monitoring network traffic and analyzing patterns using generative models, acting as an early warning system for potential zero-day attacks.
Furthermore, the effectiveness and precision of IDS can be increased using GAI approaches. IDS can improve their detection accuracy and minimize the impact of false alarms by lowering false positives and false negatives [109]. Additionally, the use of generative models can lower the computing burden of IDS, enabling instantaneous network traffic monitoring and analysis [110]. However, integrating GAI into IDSs can be difficult. The necessity for extensive wide-ranging training datasets, the interpretability of generative models, and the adversarial robustness of the models against complex attacks are a few complications.
@@ IV-C Jamming Attacks
The integrity and availability of mobile networks are seriously threatened by jamming assaults in the area of wireless security. These assaults encompass deliberate disruption or interference of wireless signals, which results in a reduction in service or prevention of communication [111]. Innovative methods for countering and lessening the effects of jamming attacks in mobile networking are possible using GAI techniques. GANs can be useful in recognizing and preventing jamming attacks. These models can recognize irregularities brought on by jamming activities because they can learn and comprehend the typical patterns and properties of wireless signals [112]. GAI models can identify the presence of jamming signals by continuously observing and analyzing network traffic and can then activate the necessary countermeasures.
To create synthetic samples that closely match real signals, one method is to train GAI models on massive datasets of typical wireless signal patterns. Any appreciable differences between the generated samples and the received signals can be used to identify potential jamming assaults [40]. This improves the responsiveness of security systems by enabling the quick and precise detection of jamming actions. Intelligent jamming mitigation solutions can be developed with the help of GAI. Models can learn to identify and categorize many sorts of jamming attacks by being trained on a wide variety of signals [14]. Using this knowledge, adaptive and robust defenses can be created to dynamically modify their tactics to overcome particular jamming approaches.
Furthermore, jamming attack prediction and prevention can be accomplished through GAI approaches. These algorithms can identify potential weaknesses and forecast the risk of upcoming jamming occurrences by examining past data and patterns [40]. This proactive strategy enables network operators to implement preventive actions to lessen the effects of prospective jamming assaults, such as modifying frequencies or adding security measures. The use of GAI for jamming attack detection and mitigation is not without challenges. These include the need for substantial diverse training datasets, the ability of models to adapt to evolving jamming strategies, and the need for real-time processing for quick response [112]. These issues can be resolved and the defense against jamming assaults in mobile networking can be strengthened with additional studies and developments in GAI.
@@ IV-D Generative Steganography
Data privacy and sensitive information security are of the highest priority for modern wireless networks. To improve data security in mobile networking, generative steganography combines the ideas of GAI and steganography to conceal sensitive information within apparently harmless wireless signals [113]. With the help of steganography, we can hide information such that undesirable recipients cannot detect it. GANs can be used to produce cover signals with secret information. These models study the statistical properties of typical wireless signals and employ this information to include encrypted data within them [114].
@@
In terms of mobile networking security, generative steganography algorithms offer several benefits. They provide a covert way to send sensitive data without prompting suspicion [101]. Generative steganography overcomes usual detection techniques by hiding data within wireless signals that appear to be harmless. Additionally, generative steganography improves mobile networks’ resistance to unauthorized access and data breaches. Since the secret data are integrated into the cover signals, it is safe and impractical to decrypt without proper tools even if attackers intercept the transmission [115].
Furthermore, generative steganography offers a strong defense against assaults meant to compromise the secrecy and authenticity of wireless communications. The process of embedding and extracting concealed information can be extremely efficient and resilient to detection by unauthorized parties using GAI models [116]. However, some issues must be resolved for generative steganography to be used in real-world settings. The challenges include maximizing the capacity of concealed information while minimizing the influence on the cover signals through optimization of embedding algorithms and the technique’s resilience to various sorts of attacks.
Figure 5: Layer-wise representation of a GAI-based IDS for privacy preservation in mobile networking.
Table IV: GAI for Optimizing Wireless Security Enforcement Scenarios.
@@
Apps Ref Model Inputs Outputs Loss function Performance
@@
Sensitive Data Obscurity [117] GAP-GAN Spectrum sharing policies Obfuscated data for adversary’s inference Minimax Balances attacks and utility for spectrum sharing and signal mapping
@@
[118] Federated DL Decentralized traffic statistics Botnet attack detection and anomaly detection results MSE Outperforms standard centrally managed system in attack detection accuracy
@@
[119] Stacking ensemble Darknet traffic data Darknet traffic identification and characterization MSE Robust against four adversarial attacks
@@
Intrusion Detection [120] CVAE Intrusion data Intrusion detection results log-cosh loss term Capable of generating new intrusion data with promising diversity
@@
[121] Improved CVAE Network data features Intrusion detection results Max Better overall accuracy, detection rate, and false positive rate
@@
[106] Autoencoder Heterogeneous network data Intrusion detection results Wasserstein distance Efficiently detects network threats and resolves data imbalance problems
@@
[122] Deep autoencoder Metaverse network traffic data IoT intrusion detection Wasserstein distance Higher accuracy for binary and multiple classification
@@
[105] Transformer Data from 6G-enabled IoT Cyber attack detection results Log discriminator Achieved high overall accuracy in detecting IoT attacks
@@
Jamming Attacks [123] GDNN, GALNN Communication data from cognitive radios Transmission decisions to mitigate jamming attacks Cross entropy Achieves successful transmission by taking optimal actions for defense mechanism
[124] SamBee Frames under different frequencies Jammed ZigBee frames in distinct channels MSE Addresses parallel spoofing attacks and multiple-channels jamming
@@
[125] GAN-enhanced Spectrum sensing information Anti-jamming channel access policy Adversarial & reconstruction Effectively avoids complex jamming attacks
@@
[126] GAN-based Original wireless signal features Probability of misclassification for spoofing attacks Minimax Significant increase in the success probability of wireless signal spoofing
@@
[127] Hidden Markov model Signals processed with sliding window Jamming status of each received energy sample Maximum posterior probability Demonstrates superiority in detecting stealthy jamming without prior knowledge
@@
Generative Steganography [114] GAN based Secret information with noise vector Stegno image containing hidden secret information Minimax Evaluation shows improved image quality and security
@@
[128] GSN Secret and noise data Realistic stego images with secret data Minimax, adversarial loss Superior resistance to steganalysis detection
@@
[34] DCGAN Secret message, uncorrupted region in the image Stego image with hidden secret message Contextual & perceptual Qualitative and quantitative evaluations of the generated stego images.
[129] LSTM Sentence for analysis Prediction of the anticipated texts Cross entropy The optimization algorithm reduces losses in neural networks
@@
[116] GAN, CNN Dealer’s distribution and human face images Revealed message when authorized participants grant consent Noise vector Secure message authentication and multi-factor authentication
IV-E Lessons Learned and Keys Takeaway

### v2

@@ III-C Channel Estimation
Channel estimation is a process used in wireless networking to determine the characteristics of a communication channel. This information is crucial for optimizing the transmission and reception of data. Numerous benefits can be achieved by incorporating GAI into mobile networking channel estimation. First, GAI enhances the ability to detect faulty signals and provide appropriate fixes to improve channel estimation procedures. Second, GAI can streamline the resources used for channel estimation, which lowers the computational complexity and energy usage of conventional estimation techniques. Third, by learning the fundamental patterns and properties of wireless channels, GAI makes channel estimation more reliable and precise, particularly in difficult situations with little training data or time-varying channel circumstances.
@@
To verify these promising traits, the work in [107] designs a high dimensional channel estimation approach based on a deep generative network combined with a compressive sensing strategy. The experiments from [107] indicate that GAI not only uses fewer channel pilots than conventional techniques but also improves the execution time of channel estimation procedures for one-bit quantized pilot measurements. Similarly, another study for wideband channel estimation [108] has demonstrated the efficacy of GANs in dealing with channel estimation in the conditions of limited channel pilots and low transmit SNR. Specifically, it significantly reduces the number of pilots by about 70%, even at mmWave and THz frequencies, without affecting the estimation errors. On another front, the work in [109] displays the potential of deep score-based generative models to multi-input multi-output systems, with high fidelity channel estimation for channel communication of sizes 64×256 with pilot density up to 25%.
III-D Lessons Learned and Key Takeaways
Several insights in various network control and optimization fields have been gained from investigating GAI applications in network management. The dynamic nature of large-scale networks and their complexity are frequently beyond the capabilities of traditional models. However, GAI can learn complex network behaviors and automate decision-making processes with the help of VAEs and GANs. As a result, GAI-based communication networks can respond more quickly in real-time to changes in traffic, load, or performance conditions.
By anticipating traffic patterns and dynamically modifying flow rules to maximize network performance, GAI facilitates the control plane within SDN to be more responsive and agile, decreasing latency and increasing total reaction times. Besides, GAI enhances the management of network slices by modeling various user needs and network circumstances. This enables more precise and effective resource allocation across various network slices, ensuring high-quality service while reducing resource waste.
GAI has also demonstrated its effectiveness in providing more accurate forecasts of network demands for resource allocation. This allows for adaptive allocation strategies that maximize network usage and improve QoS. GAI largely simplifies resource allocation procedures through reinforcement learning-based techniques, especially when network demands fluctuate. GAI’s ability to create models from sparse or insufficient data benefits channel estimation. It is therefore possible to forecast channel conditions more precisely, which enhances communication reliability—particularly in challenging situations where conventional approaches would be insufficient.
The key takeaways from this section are as follows.
@@ III-D Lessons Learned and Key Takeaways
  • Key 1. GAI can enhance network control by employing predictive models to optimize traffic flow and minimize time (e.g., training time and execution time).
  • Key 2. GAI offers flexible network routing by extracting information learned from synthetic simulation scenarios with combination of real-time monitoring of data pathways.
@@
  • Key 3. By dynamically adjusting computing resources based on real-time demand and ensuring optimal efficiency, the resource allocation capabilities are improved.
  • Key 4. Generative models improve estimation accuracy, resulting in significant benefits for channel estimation.
@@ IV Application of GAI in Wireless Security
This section reviews the applications of GAI in dealing with security issues of mobile and wireless networking, covering from sensitive data obscurity, intrusion detection, and jamming attacks, to generative steganography (See Fig. 6). Then, this section concludes with useful lessons and key takeaways.
Figure 6: Illustration of GAI applications in wireless security: sensitive data obscurity, intrusion detection system, jamming attack, and generative steganography.
@@ IV-A Sensitive Data Obscurity
Sensitive data obscurity refers to techniques that hide or obscure sensitive information to protect it from unauthorized access or eavesdropping, such as encryption and data masking [110]. This ensures that data remains unreadable and secure even if it is intercepted. In the context of large-scale IoT particularly in mobile and wireless networking, sharing raw data between intelligent systems for training machine learning models without proper safeguarding strategies poses numerous risks to user privacy, where sensitive information may be disclosed by third parties through unauthorized access or interception, as seen in the case of sensors in IoT devices [111]. Recently, GAI has demonstrated its promising capabilities in obfuscating sensitive data by creating synthetic data that closely mimics real datasets [112].
There are multiple methods for generating synthetic data using GAI models. Data perturbation is one of GAI’s main techniques in safely hiding sensitive data [113]. This technique uses synthetic data to complicate extraction of useful information, by applying controlled perturbations to sensitive data [114]. In addition, distributing and transforming only the sensitive elements in datasets can also increase the security of data sharing while preserving the features and statistical utility of the original data for model training activities [115]. Moreover, GAI models can apply other techniques like differential privacy and data anonymization. For example, in encrypted communication, indistinguishable noise can be created from encrypted data using generative models, complicating decryption of the original data without the proper decryption keys [116]. With this strategy, sensitive data are sent more securely over wireless networks while maintaining their confidentiality and integrity.
@@ IV-B Intrusion Detection
Intrusion detection is a process of monitoring and analyzing events in wireless-secured IoT networks to identify signs of unauthorized access or malicious activity. IDS is a valuable tool for detecting these threats by monitoring network traffic and raising system alarms when suspicious activity is identified. Due to its predominant features, applying GAI to IDSs opens new avenues to enhance detection capabilities and improve wireless network security [117]. GAN is a valuable tool for discovering anomalies and distinguishing malicious activity in wireless network traffic assisted by GAI models. GANs are capable of learning the underlying patterns and behaviors of network traffic through training on large-scale IoT datasets, allowing them to detect unforeseen attack patterns from expected norms with high accuracy [118].
Generating synthetic data samples that closely resemble actual network traffic is one of the main benefits of utilizing GAI in intrusion detection. IDS is more resilient and able to handle different types of threats when it uses GAIs to improve its training data and create a range of testing and evaluation scenarios based on fake attack patterns. The early discovery of zero-day attacks, vulnerabilities or exploits of which security experts are unaware, can also be helped by GAI [119], where IDS exploits generative models to detect suspicious actions that depart from usual behavior and generates an early warning action to the system for potential zero-day attacks. Additionally, IDSs-based GAI approaches can improve their detection accuracy and minimize the impact of false alarms by lowering false positives and false negatives [120]. Similarly, the use of generative models can lower the computing burden of IDS, enabling instantaneous network traffic monitoring and analysis [121].
@@ IV-C Jamming Attacks
Jamming attacks refer to the intentional transmission of interfering signals by attackers to disrupt or block wireless communication, hindering legitimate devices to communicate due to poor reception [122]. In the context of IoT where multiple networks coexist, exploiting GAI can help recognize and prevent jamming attacks by taking necessary actions when they occur, for its ability to learn and comprehend the underlying patterns and characteristics of wireless signals.
For example, GAI can help mitigate jamming action by making the transmitter’s behaviour unpredictable [123], such as taking deliberately wrong transmissions in some selected time slots (i.e., transmitting on a busy/idle channel) in combination with the ML algorithm for spectrum sensing classification. This results in an attack in which the transmitter returns to the jammer, deceiving the adversary’s surveillance capabilities. Besides, exploiting GAI brings considerable benefits to secure cognitive radio networks by extracting multi-signal representation from a wideband spectrum with a high sampling rate and consequent high-dimensional data [124]. Likewise, exploiting GAI can help the receiver classify accurate timing and positioning information in the navigation task from distinguishing between spoofed and authentic satellite signals [125]. Another example is GAI-assisted jamming mitigation systems trained on diverse signals, which can identify and categorize various types of jamming attacks [16].
@@ IV-D Generative Steganography
Generative steganography is a data-hiding method, where secret data is directly used to generate stego media, such as images, without requiring a cover image [126]. In light of wirelessly intelligent IoT networks, where data privacy and sensitive information security are top priorities, the combination of GAI and steganography enhances safeguarding sensitive information in seemingly harmless wireless signals [127]. In this approach, GAI uses GANs, trained on the statistical properties of typical wireless signals [128], to generate realistic images and embed secure information directly into the pixel values, and the steganography techniques provide methods to securely embed and extract this hidden data.
@@
In this way, bringing GAI-combined steganography into mobile and wireless networking offers several benefits. For example, this approach provides a covert way to send sensitive data without prompting suspicion [112]. Besides, GAI-combined steganography improves mobile networks’ resistance to unauthorized access and data breaches [129]. Even if the attackers can successfully intercept the transmission, decrypting the secret data integrated into the cover signals without proper tools is impractical. Furthermore, generative steganography offers a strong defense against assaults meant to compromise the secrecy and authenticity of wireless communications. The process of embedding and extracting concealed information can be extremely efficient and resilient to detection by unauthorized parties using GAI models [130].
Table IV: GAI Use Cases in Wireless Network Security
Security Applications Techniques Problems Addressed Observed Benefits Tangible Impact
Sensitive data obscurity [111] VAEs Safeguarding private information Improved anonymization of data Enhanced adherence to privacy laws
Intrusion detection [117] GANs Identifying any unapproved access Enhanced precision in identifying breaches Increased robustness of the network
Jamming attacks [124] Diffusion Models Reducing interference caused by jammers Real-time jamming attack mitigation Better accessibility of the network
Generative steganography [131] GANs Data hiding mechanisms Improved data protection and camouflage Routes of secure communication
Secure data transmission [31] Autoencoders Hiding sensitive data during transmission Improved security of data in transit Decreased chance of data leaks
Traffic anomaly detection [42] GANs Recognizing unusual traffic trends Enhanced identification of possible threats Enhanced security protocols for networks
Physical layer security [29] Diffusion Models Addressing the physical layer’s vulnerabilities Increased channel security Enhanced honesty and secrecy
Table V: GAI for Optimizing Wireless Security Enforcement Scenarios.
@@
Apps Ref Model Inputs Outputs Loss function Performance Limitations
@@
Sensitive Data Obscurity [132] GAP-GAN Spectrum sharing policies Obfuscated data for adversary’s inference Minimax Balances attacks and utility for spectrum sharing and signal mapping GAN architecture may lead to mode collapse, where the privatizer could dominate the adversary.
@@
[133] Federated DL Decentralized traffic statistics Botnet attack detection and anomaly detection results MSE Outperforms standard centrally managed system in attack detection accuracy Models may suffer from data heterogeneity.
@@
[134] Stacking ensemble Darknet traffic data Darknet traffic identification and characterization MSE Robust against four adversarial attacks The learners are simple; thus, not robust on complicated data.
@@
Intrusion Detection [135] CVAE Intrusion data Intrusion detection results log-cosh loss term Capable of generating new intrusion data with promising diversity Requirements of labels make the method infeasible when applied to real data that lacks human labellings.
@@
[136] CVAE Network data features Intrusion detection results Max Better overall accuracy, detection rate, and false positive rate Requirements of label make the method infeasible when applied to real data that lacks human labellings.
@@
[117] GAN Heterogeneous network data Intrusion detection results GAN Loss Efficiently detects network threats and resolves data imbalance problems GAN architecture may lead to mode collapse, where the privatizer could dominate the adversary.
@@
[137] GAN Metaverse network traffic data IoT intrusion detection GAN Loss Higher accuracy for binary and multiple classification GAN architecture may lead to mode collapse where the privatizer could dominate the adversary.
@@
[116] Transformer-based GAN Data from 6G-enabled IoT Cyber attack detection results GAN Loss Achieved high overall accuracy in detecting IoT attacks GAN architecture may lead to mode collapse, where the privatizer could dominate the adversary.
@@
Jamming Attacks [138] GDNN, GALNN Communication data from cognitive radios Transmission decisions to mitigate jamming attacks Cross entropy Achieves successful transmission by taking optimal actions for defense mechanism GAN architecture may lead to mode collapse, where the privatizer could dominate the adversary.
@@
[139] GAN-enhanced Spectrum sensing information Reconstructed spectrum waterfall Adversarial & reconstruction Effectively avoids complex jamming attacks The need for data accumulation in the experience memory causes the system to suffer from straggling.
@@
[140] GAN-based Modulation signal Probability of misclassification for spoofing attacks Minimax Significant increase in the success probability of wireless signal spoofing The model is fixed on the trained modulation signals, and may not generalize to other signals.
@@
[141] Hidden Markov model Signals processed with sliding window Jamming status of each received energy sample MSE Demonstrates superiority in detecting stealthy jamming without prior knowledge The simple model makes the learning not generalized when detecting the jammers with different behaviors.
@@
Generative Steganography [128] GAN based Secret information with noise vector Stegno image containing hidden secret information Minimax Evaluation shows improved image quality and security A huge amount of data is required to train the GAN model.
@@
[142] GSN Secret and noise data Realistic stego images with secret data GAN loss Superior resistance to steganalysis detection The model may suffer from mode collapse, where the data is overfitted into another data.
@@
[31] DCGAN Secret message, uncorrupted region in the image Stego image with hidden secret message Contextual & perceptual Qualitative and quantitative evaluations of the generated stego images. The model may suffer from mode collapse, where the data is overfitted into another data.
@@
[130] GAN Dealer’s distribution and human face images Revealed message when authorized participants grant consent Noise vector Secure message authentication and multi-factor authentication The training process needs high computational resources, and the model is large, making it infeasible for IoT
IV-E Lessons Learned and Key Takeaways